A solution of washing soda ( sodium carbonate, Na2CO3 ) has a hydroxide ion concentration of 0.004 mol dm-3. What is the pH at 25C ?


1
Expert's answer
2021-03-02T02:01:06-0500

pH = 14 - pOH = 14 + lg(0.004) = 11.6